OpenStudy (anonymous):

Find the roots of Z^(2) + (2-3I)Z -6i

OpenStudy (anonymous):

Use \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\] where \[a =1\] \[b =2-3i\] \[c =-6i\]

OpenStudy (anonymous):

U ll get \[z =\frac{ -(2-3i)\pm \sqrt{-5+12i} }{ 2 }\]

OpenStudy (zehanz):

It is probably not enough to just fill in the abc-formula. Mostly it is required to give the answer in a+bi form. Now sqrt(-5+12i) = 2+3i, so\[z_1=(-2+3i+2+3i)/2 =3i\]and\[z_2=(-2+3i-2-3i)/2=-4/2=-2\]

OpenStudy (anonymous):

@ZeHanz how did you go from sqrt(-5+12i) to 2+3i

OpenStudy (anonymous):

@ZeHanz Can you answer my question please i am really confused =D

OpenStudy (zehanz):

Sorry about that! I admit I use my calculator sometimes for these square roots...but when it gives you a "nice" number, it is worth while trying to get it without the help of a calculator. Here is how you could do that: we look for a number a+bi, such that (a+bi)²=-5+12i. After expanding the brackets, you get a²-b²+2abi=-5+12i, so: a²-b²=-5 and 2ab=12, so ab=6. ab=6 doesn't leave may options open for "nice" numbers, so why not try a=2 and b=3? Yes! they do the trick! 2²-3²=4-9=-5. I wouldn't know another way of doing it, besides halving the argument and taking the square root of the magnitude. In this case that doesnt work well, because the argument of -5+12i is itself not a "nice" angle. I hope your a little less confused now!